• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name X. Zheng

    Log entry time 16:07:14 on February17,2009

    Entry number 261649

    keyword=Shift Summary

    8:00 Shift started with Hall in controlled access and accelerator people fixing a Moller power supply. Beam study are starting and should last until 2pm.

    8:15 Jack is accessing the Hall to ramp down Q3 and possibly the dipole for left HRS.

    8:45 Jack called saying that the Moller PS are run at too high currents. They smell and this triggered the VESDA system. We called MCC but they don't know the status of the PS work. They gave us a pager to call to talk to the people working on the moller PS in the service building. Jack will call. Meanwhile, Ed called them already and asked them to shut down the PS.

    13:50 Everyone is out of the hall, going back to beam permit.

    14:38 Set BB Shower T6 threshold to -120mV.

    15:00 MCC has problem setting up the current to 30uA

    16:00 MCC can deliver 30uA and will call us when they are ready for BCM calibration, continue to next shift